OpenStudy (anonymous):

HELP PLEASE! How many mL are in 27 hL?

OpenStudy (danjs):



OpenStudy (danjs):

Only think you need to remember , is how each prefix relates to the base unit, Liter

OpenStudy (anonymous):

ohhh okay, so they all cancel out XD

OpenStudy (anonymous):

i just have one question

OpenStudy (danjs):

right, you start with what you are gven 27 hL, then multiply by ratios of 1, to end with the units you want

OpenStudy (danjs):

27 x 10^5 mL

OpenStudy (anonymous):

so i would multiply the numbers on the top?

OpenStudy (danjs):

The way i always did these types of probs, and stoichiometry probs, just start with what you are given, dont worry about the numbers, then arrange the ratios of units to cancel correctly, at the very end, just multiply through all the numbers in the top, and divide by all the numbers in the bottom

OpenStudy (danjs):

yeah, vertical line = multiply, horizontal line = divide

OpenStudy (anonymous):

oh yes, that seems like a good idea :P sometimes i get confused cause some units don't cancel out correctly.

OpenStudy (danjs):

Just do them backwards to what you are used to, the units come first, then worry about the #s

OpenStudy (danjs):

if you want to post another prob for practice, i can help you though it

OpenStudy (anonymous):

i have a question that may seem kinda dumb lol, nut here it goes: when it says divide by all the numbers on the bottom for this problem, do i literally divide it by 10^3?

OpenStudy (danjs):

10^(-3) L = 1 ml divide by 10^(-3) is the same as multiplying by 10^(+3) negative exponents you can move to the other side of the fraction and make positive

OpenStudy (danjs):

or if you have a decent calculator, you can type in divide by 10^(-3) if you want

OpenStudy (danjs):

Multiply across all the top numbers,, Then divide by all the bottom numbers

OpenStudy (anonymous):

lol, thats the prob, i don't have a calculator for this, thats why im having so much trouble :P

OpenStudy (danjs):

right , so 10^(2) divided by 10^(-3) is the same as \[\frac{ 10^2 }{ 10^{-3} } = \frac{ 10^2 *10^{3} }{ 1 } = \frac{ 10^{3+2} }{ 1 } = 10^5\]

OpenStudy (danjs):

so the answer is just 27 x 10^5

OpenStudy (danjs):

or 2700000 mL

OpenStudy (anonymous):

Ohhhh I finally see it now!!! XD Omg Thanks!

OpenStudy (danjs):

no, prob post more if you want to run through a couple

OpenStudy (danjs):

those are all you need to remember, how each prefix relates to the base unit Liter, Meter, Gram

OpenStudy (anonymous):

Thank you. It seems so much clearer now, I was really struggling to understand anything before :P

OpenStudy (danjs):

cool, post another if you want to practice more

OpenStudy (anonymous):

okay :D

OpenStudy (danjs):

new question thread

OpenStudy (anonymous):

okay